OpenStudy (anonymous):

Okay: square root of -196 im about to have a test over this so imma need somebody to dumb it down

OpenStudy (lgbasallote):

\[\Large \sqrt{-196} \implies \sqrt{196 \times -1} \implies \sqrt{196} \times \sqrt{-1}\] does that help?

OpenStudy (anonymous):

first of all it must be a complex number, because there is no real square root of a negative number. is that ok?

OpenStudy (anonymous):

what the flip where did you get a one there is no one dumb butt

OpenStudy (anonymous):

secondly oh nmv

OpenStudy (phi):

@delaine_22 is it necessary to be rude? People are both trying to help and are correct.

OpenStudy (anonymous):

nobody asked you to but in the conversation anyway and yeah i was askinga question so yeah if theyu think im being rude ohh welll idc i was asking a question so if you dont like the way i ask questions dont look at my stuff <3 thanks (: